We assume given a plane U passing by the tip of the vectors ~u =< −1, 1, 2 >, ~v =< 2, −1, 0 > and w~ =< 1, 1, 3 >.

(6.1) Find the dot products ~u · ~v and w~ · ~v

(6.2) Determine whether or not there is a vector ~n that is perpendicular to U. If yes, then find the vector ~n. Otherwise explain why such a vector does not exist?


1
Expert's answer
2021-06-08T12:57:44-0400

(6.1)


"\vec u\cdot\vec v=-1(2)+1(-1)+2(0)=-3"

"\vec w\cdot\vec v=1(2)+1(-1)+3(0)=1"

(6.2)


"\vec u\times\vec v=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 -1 & 1 & 2 \\
 2 & -1 & 0 \\
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 1 & 2 \\
 -1 & 0
\end{vmatrix}-\vec j\begin{vmatrix}
 -1 & 2 \\
 2 & 0
\end{vmatrix}+\vec k\begin{vmatrix}
 -1 & 1 \\
 2 & -1
\end{vmatrix}"

"=2\vec i+4\vec j-\vec k"



"\vec u\times\vec w=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 -1 & 1 & 2 \\
 1 & 1 & 3 \\
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 1 & 2 \\
 1 & 3
\end{vmatrix}-\vec j\begin{vmatrix}
 -1 & 2 \\
 1 & 3
\end{vmatrix}+\vec k\begin{vmatrix}
 -1 & 1 \\
 1 & 1
\end{vmatrix}"

"=\vec i+5\vec j-2\vec k"


Two vectors "2\vec i+4\vec j-\vec k" and "\vec i+5\vec j-2\vec k" are not collinear.

Three vectors are not in the single plane. Then we cannot find the vector "\vec n." Such vector "\vec n" does not exist.